OpenStudy (anonymous):

an octagon has side lengths equal to 16.35 in. Find the area.

OpenStudy (anonymous):

Area of a polygon of n sides and side length b is \[ \frac { n b^2}4 \cot (\pi/n)=\frac{1}{4} \left(8( 16.35^2)\right) \cot \left(\frac{\pi }{8.}\right)= 1290.75 \]